# Training Versatile Coding Agents in Synthetic Environments

### v1

@@ 7 Conclusion and Future Work
While these results are promising, we identify two key directions for future work. Firstly, incorporating a broader set of benchmarks, such as SWE-bench Multimodal and SWE-perf, would further validate the extensibility of our framework, as their features align well with the generative capabilities of SWE-Playground pipeline. Secondly, conducting RL experiments is an interesting direction to explore the full potential of our environment generation pipeline. As we have argued, our pipeline creates a scenario where unit test generation and code implementation mutually verify each other. Training an agent to master both skills through RL could directly probe its capacity to develop self-verifying and, ultimately, self-improving coding agents like liu2025agent0vl.
@@ Appendix A Related Work
Coding Agents. Capable Large Language Models and robust agent scaffolds are two critical components of modern coding agents. Recent state-of-the-art language models include proprietary ones such as Claude Sonnet 4.5 (anthropic2025claude45) and GPT-5 (openai2025gpt5), as well as open-source models such as Qwen3-Coder (qwen3technicalreport), Kimi K2 (kimiteam2025kimik2openagentic) and DeepSeek-V3.1 (deepseekai2024deepseekv3technicalreport). For coding agents scaffolds, SWE-agent (yang2024sweagent) introduces the concept of agent-computer interface which offers a small set of simple actions for interacting with files. Building upon SWE-agent, mini-SWE-agent (yang2024sweagent) achieves similar performance with a simpler control flow. OpenHands (wang2025openhands) expands on this by offering a wider action space including web browsing and API calls, enabling models to tackle more complex and diverse tasks. SWE-search  (antoniades2024swesearchenhancingsoftwareagents) employs a multi-agent system and Monte Carlo Tree Search (MCTS) strategy for planning to enhance flexibility during state transition. The combination of these powerful models and agent scaffolds enables coding agents to solve real-world coding challenges and even surpass the performance of human engineers on certain tasks (li2025aiteammates) .
@@
Coding Tasks and Benchmarks. HumanEval (chen2021codex) is the first coding benchmark to use an executable environment for verification instead of relying on vague text similarity, laying the foundation for subsequent work. LiveCodeBench (jain2024livecodebench) advances this by increasing the difficulty and evaluation coverage, gathering problems from contests across various competition platforms. BigCodeBench (zhuo2024bigcodebench) challenges models to follow complex instructions to invoke function calls as tools for solving practical tasks. In addition to these function-level and file-level edition tasks, a large number of recent benchmarks have been designed to test the performance of models and agents on repository-level coding tasks. Commit-0 (zhao2024commit0librarygenerationscratch) targets rebuilding existing libraries from scratch. SWE-bench (jimenez2024swebench) is also crafted from real GitHub repositories and challenges models to solve issues noted in pull requests. SWE-Bench Pro (deng2025swebenchproaiagents) builds upon SWE-bench with more challenging tasks to capture realistic, complex, enterprise-level problems beyond the scope of SWE-Bench. SWE-bench Multimodal (yang2024swebenchmultimodal) and SWE-bench Multilingual (yang2025swesmith) both extend SWE-bench to evaluate more diverse abilities. SWE-bench Multimodal incorporates visual elements such as diagrams and screenshots to test whether modals can perform well on visual information. SWE-bench Multilingual covers nine different programming languages for the evaluation of various application domains. Multi-SWE-bench (zan2025multiswebench) is also a issue-resolving benchmark aiming to evaluate various programming languages. SWT-Bench (mundler2024swtbench) evaluates model proficiency on reproducing GitHub issues in Python. TestGenEval (jain2025testgenevalrealworldunit) measures test generation performance in well-maintained Python repositories. ArtifactsBench (zhang2025artifactsbenchbridgingvisualinteractivegap) assesses the ability to transform multimodal instructions into high-quality and interactive visual artifacts. SWE-Perf (he2025sweperf) and SWE-fficiency (ma2025swefficiencylanguagemodelsoptimize) benchmarks models on optimizing code performance in real-world repositories. KernelBench (ouyang2025kernelbenchllmswriteefficient) evaluates model capability on writing efficient GPU kernels. We argue that all the evaluated abilities are equally important for coding agents, and the exceptional performance on a single benchmark is not indicative of a truly capable agent.
@@
HumanEval (chen2021codex) 2021.70 164 Static Fuction-Level Functional Correctness Human Crafted
LiveCodeBench (jain2024livecodebench) 2024.30 511 Static Holistic Coding Abilities on Contest Problems Auto Collected
BigCodeBench (zhuo2024bigcodebench) 2024.60 1,140 Static Complex Instruction Following with API Calling Human Crafted
Commit-0 (zhao2024commit0librarygenerationscratch) 2024.90 54 Agentic Repository-Level Library Generation From Scratch Auto Collected
SWE-bench (jimenez2024swebench) 2023.10 2,294 Agentic GitHub Issue Resolution Auto Collected
SWE-bench Pro (deng2025swebenchproaiagents) 2025.90 1,865 Agentic Long-Horizon Enterprise-Level Multi-File Patches Mixed Curation
SWE-bench Multimodal (yang2024swebenchmultimodal) 2024.10 517 Agentic Visual Context Understanding & Generation Auto Collected
SWE-bench Multilingual (yang2025swesmith) 2025.50 300 Agentic Multiple Programming Languages Auto Collected
Multi-SWE-bench (zan2025multiswebench) 2025.40 4,723 Agentic Multilingual Issue Resolution Auto Collected
SWT-Bench (mundler2024swtbench) 2024.60 1,900 Agentic Test Generation for Bug Reproduction Auto Collected
TestGenEval (jain2025testgenevalrealworldunit) 2024.10 1,210 Static Paired Test Generation & Completion Human Crafted
ArtifactsBench (zhang2025artifactsbenchbridgingvisualinteractivegap) 2025.70 1,825 Static Interactive Visual Artifacts Human Crafted
SWE-Perf (he2025sweperf) 2025.70 140 Agentic Code Performance Optimization Auto Collected
SWE-fficiency (ma2025swefficiencylanguagemodelsoptimize) 2025.11 498 Agentic Repository-Level Performance Optimization Human Crafted
KernelBench (ouyang2025kernelbenchllmswriteefficient) 2025.20 250 Static Efficient PyTorch Kernel Generation Human Crafted
@@ B.1 Training Details
Agent Scaffold. Following pan2025training and jain2025r2e, we use OpenHands (wang2025openhands) CodeAct Agent as our agent scaffold. OpenHands CodeAct Agent leverages ReAct framework (yao2023react), empowering agents with the ability to interact with the command line, edit files and browse the web based on a reasoning and acting loop. OpenHands also provides securely isolated Docker sandboxes for all the actions to be executed, thereby ensuring a safe and consistent execution environment that prevents unintended changes to the host system. In our experiments, we use version 0.48.0222https://github.com/All-Hands-AI/OpenHands/releases/tag/0.48.0 of OpenHands and disable the browsing tool and function call features.
@@
Dataset Collection and Statistics. We employ an ensemble of Gemini 2.5 Pro (comanici2025gemini25pushingfrontier), GPT-4.1 (openai2025gpt41) and Claude Sonnet 4 (anthropic2025claude4) for proposing projects. For proposing step-by-step tasks, we use GPT-4.1. All agents involved in our pipeline are implemented using OpenHands in conjunction with Claude Sonnet 4. We collect from 28 generated projects a total of 280 general trajectories (as detailed in Section 3.1 to Section 3.5) and 424 capability-specific trajectories (as detailed in Section 4), respectively 213 for issue resolution, 183 for issue reproduction and 28 for library generation, with Claude Sonnet 4. In experiments, we exclusively use the trajectories of finishing the task instead of those for generating the task. While the trajectories for unit test generation and issue injection could be valuable for training agents on specialized tasks such as generating unit tests from documentation, we do not include such benchmarks in our experiments, with the exception of SWT-Bench (mundler2024swtbench), which requires agents to reproduce issues, but not to purely write unit tests. Therefore, we only adopt the trajectories that were are relevant to the capabilities evaluated in the benchmarks, namely, the trajectories of building up the project, fixing bugs and reproducing issues.
@@
Training details. Following previous works (pan2025training; jain2025r2e; yang2025swesmith), we select Qwen-2.5-Coder series (hui2024qwen2), Qwen-2.5-Coder-7B-Instruct and Qwen-2.5-Coder-32B-Instruct, as our base models. We finetune the base model with all the trajectories mixed together and randomly shuffled, using a learning rate of 1e-5 and training for 3 epochs. Given that the length of our functionality implementation trajectories frequently exceeds the default maximum context length of 32,768 tokens, we follow the official instructions [^3] https://huggingface.co/Qwen/Qwen2.5-Coder-7B-Instruct to extend context length using YaRN (peng2024yarn) during both training and inference stage. Following jain2025r2e, we use LLaMA-Factory  (zheng2024llamafactory) for efficient model finetuning, and adopt zou360lf to further integrate sequence parallelism feature.
@@
Benchmarks and Baselines. We select three distinct benchmarks to demonstrate the effectiveness of SWE-Playground: SWE-bench, SWT-Bench and Commit-0. SWE-bench (jimenez2024swebench) focuses on solving real GitHub issues. SWT-Bench (mundler2024swtbench) evaluates agents on their ability to reproduce issues. Commit-0 (zhao2024commit0librarygenerationscratch) tasks agents with building Python libraries from scratch. Collectively, these three benchmarks cover a wide range of critical coding capabilities, allowing for a comprehensive understanding on whether agents can handle various tasks. We adopt SWE-bench Verified (chowdhury2024swebenchverified), SWT-Bench Lite and Commit-0 Lite as the evaluation splits for the respective benchmarks. We report resolved rate and empty patch rate for SWE-bench Verified, resolved rate and coverage delta for SWT-Bench Lite, and resolved rate for Commit-0 Lite. We select three strong baselines that focus on constructing coding environments and training SWE agents: SWE-Gym (pan2025training), R2E-Gym (jain2025r2e) and SWE-smith (yang2025swesmith). Similar to our approach, SWE-Gym and R2E-Gym also build upon OpenHands, while SWE-smith is based on SWE-agent framework (yang2024sweagent). We use vLLM (kwon2023efficient) to serve all the models for inference. We use OpenHands remote runtime (neubig2024-evaluation-of-llms) to host docker containers during evaluation.
@@ Appendix C Extended Results
We provide our full experimental results in Table 6. In addition to SWE-bench, SWT-Bench and Commit-0, we also evaluate models on BigCodeBench (zhuo2024bigcodebench), a non-agentic benchmark which assesses models on programming challenges. As the results show, all evaluated models demonstrate performance degradation against the base model on this benchmark, with SWE-smith-7B performing the worst with only a $6.08\%$ success rate. This result validates that training with agentic trajectories inevitably comes at the cost of the capability on non-agentic static tasks. Firstly, models trained with agentic tasks typically adapt to the complex interactions and execution-based verification nature of agentic environments, and thus falter when facing static coding tasks. Secondly, agentic tasks such as SWE-bench are inherently different in format and capabilities evaluated from non-agentic tasks such as BigCodeBench, with the former focusing on context and project understanding, while the latter stressing algorithm design and instruction following. Furthermore, we hypothesize that this effect might be largely influenced by the design of the agent scaffold used during training.
@@
Qwen2.5-Coder-7B - 01.81.8 $45.8$ 00.720.72 07.557.55 05.215.21 $17.58$
SWE-Gym-7B 491 $10.6$ $33.8$ 01.451.45 05.195.19 07.197.19 $12.16$
R2E-Gym-7B 3.3k $19.0$ 00.720.72 02.662.66 07.287.28 $16.90$
SWE-smith-7B 2.0k $15.2$ 00.000.00 $12.30$ 07.507.50 06.086.08
SWE-Play-mix-7B 704 $17.0$ 06.46.4 04.484.48 $24.15$ 08.468.46 $14.19$
SWE-Play-general-7B 280 08.68.6 $22.6$ 02.542.54 07.557.55 07.507.50 $14.86$
SWE-Play-swe-7B 213 $11.0$ $10.6$ 00.720.72 04.804.80 06.816.81
@@
Qwen2.5-Coder-32B - 07.07.0 09.59.5 09.429.42 $24.81$ 06.626.62
SWE-Gym-32B 491 $20.6$ $13.8$ 03.263.26 $10.34$ 07.197.19
R2E-Gym-32B 3.3k $34.4$ 03.263.26 $15.04$ 09.459.45
SWE-smith-32B 5.0k $40.2$ $13.77$ $42.15$ 07.507.50
SWE-Play-mix-32B 704 $31.2$ 06.86.8 $18.12$ $44.24$ $10.42$

### v2

@@ 7 Conclusion and Future Work
While these results are promising, we identify two key directions for future work. Firstly, incorporating a broader set of benchmarks, such as SWE-bench Multimodal and SWE-perf, would further validate the extensibility of our framework, as their features align well with the generative capabilities of SWE-Playground pipeline. Secondly, conducting RL experiments is an interesting direction to explore the full potential of our environment generation pipeline. As we have argued, our pipeline creates a scenario where unit test generation and code implementation mutually verify each other. Training an agent to master both skills through RL could directly probe its capacity to develop self-verifying and, ultimately, self-improving coding agents like Liu et al. (2025).
Impact Statement
In this paper, we present SWE-Playground, an automated pipeline to synthesize environments and trajectories for the training of versatile coding agents. We anticipate several potential impacts of this work. First, SWE-Playground offers a pathway to advance the field of automated software engineering. As discussed, real-world software engineering requires a diverse set of skills beyond simple code completion, and SWE-Playground offers a robust framework for incorporating and training these versatile capabilities. From a broader perspective, with more research attention focused on the purely synthetic training data and self-evolving agents, SWE-Playground shows promising results for further research regarding these topics.
At the same time, although our experiments demonstrate the effectiveness of the SWE-Playground pipeline, rigorous quality control is essential if these systems are to be deployed in high-stakes scenarios. Furthermore, generating unit tests with LLMs remains an emerging research topic. Currently, there is a scarcity of benchmarks designed to evaluate this specific capability, and few works have discussed it systematically. Consequently, further research is required before models can be fully and reliably incorporated for automated verification.
@@ Appendix A Related Work
Coding Agents. Capable Large Language Models and robust agent scaffolds are two critical components of modern coding agents. Recent state-of-the-art language models include proprietary ones such as Claude Sonnet 4.5 (Anthropic, 2025b) and GPT-5 (OpenAI, 2025b), as well as open-source models such as Qwen3-Coder (Qwen Team, 2025), Kimi K2 (Kimi Team et al., 2025) and DeepSeek-V3.1 (DeepSeek-AI, 2024). For coding agents scaffolds, SWE-agent (Yang et al., 2024) introduces the concept of agent-computer interface which offers a small set of simple actions for interacting with files. Building upon SWE-agent, mini-SWE-agent (Yang et al., 2024) achieves similar performance with a simpler control flow. OpenHands (Wang et al., 2025b) expands on this by offering a wider action space including web browsing and API calls, enabling models to tackle more complex and diverse tasks. SWE-search  (Antoniades et al., 2024) employs a multi-agent system and Monte Carlo Tree Search (MCTS) strategy for planning to enhance flexibility during state transition. The combination of these powerful models and agent scaffolds enables coding agents to solve real-world coding challenges and even surpass the performance of human engineers on certain tasks (Li et al., 2025) .
@@
Coding Tasks and Benchmarks. HumanEval (Chen et al., 2021) is the first coding benchmark to use an executable environment for verification instead of relying on vague text similarity, laying the foundation for subsequent work. LiveCodeBench (Jain et al., 2024) advances this by increasing the difficulty and evaluation coverage, gathering problems from contests across various competition platforms. BigCodeBench (Zhuo et al., 2024) challenges models to follow complex instructions to invoke function calls as tools for solving practical tasks. In addition to these function-level and file-level edition tasks, a large number of recent benchmarks have been designed to test the performance of models and agents on repository-level coding tasks. Commit-0 (Zhao et al., 2024) targets rebuilding existing libraries from scratch. SWE-bench (Jimenez et al., 2024) is also crafted from real GitHub repositories and challenges models to solve issues noted in pull requests. SWE-Bench Pro (Deng et al., 2025) builds upon SWE-bench with more challenging tasks to capture realistic, complex, enterprise-level problems beyond the scope of SWE-Bench. SWE-bench Multimodal (Yang et al., 2025a) and SWE-bench Multilingual (Yang et al., 2025b) both extend SWE-bench to evaluate more diverse abilities. SWE-bench Multimodal incorporates visual elements such as diagrams and screenshots to test whether modals can perform well on visual information. SWE-bench Multilingual covers nine different programming languages for the evaluation of various application domains. Multi-SWE-bench (Zan et al., 2025) is also a issue-resolving benchmark aiming to evaluate various programming languages. SWT-Bench (Mündler et al., 2024) evaluates model proficiency on reproducing GitHub issues in Python. TestGenEval (Jain et al., 2025a) measures test generation performance in well-maintained Python repositories. ArtifactsBench (Zhang et al., 2025a) assesses the ability to transform multimodal instructions into high-quality and interactive visual artifacts. SWE-Perf (He et al., 2025) and SWE-fficiency (Ma et al., 2025) benchmarks models on optimizing code performance in real-world repositories. KernelBench (Ouyang et al., 2025) evaluates model capability on writing efficient GPU kernels. We argue that all the evaluated abilities are equally important for coding agents, and the exceptional performance on a single benchmark is not indicative of a truly capable agent.
@@
HumanEval (Chen et al., 2021) 2021.70 164 Static Fuction-Level Functional Correctness Human Crafted
LiveCodeBench (Jain et al., 2024) 2024.30 511 Static Holistic Coding Abilities on Contest Problems Auto Collected
BigCodeBench (Zhuo et al., 2024) 2024.60 1,140 Static Complex Instruction Following with API Calling Human Crafted
Commit-0 (Zhao et al., 2024) 2024.90 54 Agentic Repository-Level Library Generation From Scratch Auto Collected
SWE-bench (Jimenez et al., 2024) 2023.10 2,294 Agentic GitHub Issue Resolution Auto Collected
SWE-bench Pro (Deng et al., 2025) 2025.90 1,865 Agentic Long-Horizon Enterprise-Level Multi-File Patches Mixed Curation
SWE-bench Multimodal (Yang et al., 2025a) 2024.10 517 Agentic Visual Context Understanding & Generation Auto Collected
SWE-bench Multilingual (Yang et al., 2025b) 2025.50 300 Agentic Multiple Programming Languages Auto Collected
Multi-SWE-bench (Zan et al., 2025) 2025.40 4,723 Agentic Multilingual Issue Resolution Auto Collected
SWT-Bench (Mündler et al., 2024) 2024.60 1,900 Agentic Test Generation for Bug Reproduction Auto Collected
TestGenEval (Jain et al., 2025a) 2024.10 1,210 Static Paired Test Generation & Completion Human Crafted
ArtifactsBench (Zhang et al., 2025a) 2025.70 1,825 Static Interactive Visual Artifacts Human Crafted
SWE-Perf (He et al., 2025) 2025.70 140 Agentic Code Performance Optimization Auto Collected
SWE-fficiency (Ma et al., 2025) 2025.11 498 Agentic Repository-Level Performance Optimization Human Crafted
KernelBench (Ouyang et al., 2025) 2025.20 250 Static Efficient PyTorch Kernel Generation Human Crafted
@@ B.1 Training Details
Agent Scaffold. Following Pan et al. (2025) and Jain et al. (2025b), we use OpenHands (Wang et al., 2025b) CodeAct Agent as our agent scaffold. OpenHands CodeAct Agent leverages ReAct framework (Yao et al., 2023), empowering agents with the ability to interact with the command line, edit files and browse the web based on a reasoning and acting loop. OpenHands also provides securely isolated Docker sandboxes for all the actions to be executed, thereby ensuring a safe and consistent execution environment that prevents unintended changes to the host system. In our experiments, we use version 0.48.0222https://github.com/All-Hands-AI/OpenHands/releases/tag/0.48.0 of OpenHands and disable the browsing tool and function call features.
@@
Dataset Collection and Statistics. We employ an ensemble of Gemini 2.5 Pro (Comanici and others, 2025), GPT-4.1 (OpenAI, 2025a) and Claude Sonnet 4 (Anthropic, 2025a) for proposing projects. For proposing step-by-step tasks, we use GPT-4.1. All agents involved in our pipeline are implemented using OpenHands in conjunction with Claude Sonnet 4. We collect from 28 generated projects a total of 280 general trajectories (as detailed in Section 3.1 to Section 3.5) and 424 capability-specific trajectories (as detailed in Section 4), respectively 213 for issue resolution, 183 for issue reproduction and 28 for library generation, with Claude Sonnet 4. In experiments, we exclusively use the trajectories of finishing the task instead of those for generating the task. While the trajectories for unit test generation and issue injection could be valuable for training agents on specialized tasks such as generating unit tests from documentation, we do not include such benchmarks in our experiments, with the exception of SWT-Bench (Mündler et al., 2024), which requires agents to reproduce issues, but not to purely write unit tests. Therefore, we only adopt the trajectories that were are relevant to the capabilities evaluated in the benchmarks, namely, the trajectories of building up the project, fixing bugs and reproducing issues.
@@
Training details. Following previous works (Pan et al., 2025; Jain et al., 2025b; Yang et al., 2025b), we select Qwen-2.5-Coder series (Hui et al., 2024), Qwen-2.5-Coder-7B-Instruct and Qwen-2.5-Coder-32B-Instruct, as our base models. We finetune the base model with all the trajectories mixed together and randomly shuffled, using a learning rate of 1e-5 and training for 3 epochs. Given that the length of our functionality implementation trajectories frequently exceeds the default maximum context length of 32,768 tokens, we follow the official instructions [^3] https://huggingface.co/Qwen/Qwen2.5-Coder-7B-Instruct to extend context length using YaRN (Peng et al., 2024) during both training and inference stage. Following Jain et al. (2025b), we use LLaMA-Factory  (Zheng et al., 2024) for efficient model finetuning, and adopt Zou et al. (2025) to further integrate sequence parallelism feature.
@@
Benchmarks and Baselines. We select three distinct benchmarks to demonstrate the effectiveness of SWE-Playground: SWE-bench, SWT-Bench and Commit-0. SWE-bench (Jimenez et al., 2024) focuses on solving real GitHub issues. SWT-Bench (Mündler et al., 2024) evaluates agents on their ability to reproduce issues. Commit-0 (Zhao et al., 2024) tasks agents with building Python libraries from scratch. Collectively, these three benchmarks cover a wide range of critical coding capabilities, allowing for a comprehensive understanding on whether agents can handle various tasks. We adopt SWE-bench Verified (Chowdhury et al., 2024), SWT-Bench Lite and Commit-0 Lite as the evaluation splits for the respective benchmarks. We report resolved rate and empty patch rate for SWE-bench Verified, resolved rate and coverage delta for SWT-Bench Lite, and resolved rate for Commit-0 Lite. We select three strong baselines that focus on constructing coding environments and training SWE agents: SWE-Gym (Pan et al., 2025), R2E-Gym (Jain et al., 2025b) and SWE-smith (Yang et al., 2025b). Similar to our approach, SWE-Gym and R2E-Gym also build upon OpenHands, while SWE-smith is based on SWE-agent framework (Yang et al., 2024). We use vLLM (Kwon et al., 2023) to serve all the models for inference. We use OpenHands remote runtime (Neubig and Wang, 2024) to host docker containers during evaluation.
@@ Appendix C Extended Results
We provide our full experimental results in Table 6. In addition to SWE-bench, SWT-Bench and Commit-0, we also evaluate models on BigCodeBench (Zhuo et al., 2024), a non-agentic benchmark which assesses models on programming challenges. As the results show, all evaluated models demonstrate performance degradation against the base model on this benchmark, with SWE-smith-7B performing the worst with only a $6.08\%$ success rate. This result validates that training with agentic trajectories inevitably comes at the cost of the capability on non-agentic static tasks. Firstly, models trained with agentic tasks typically adapt to the complex interactions and execution-based verification nature of agentic environments, and thus falter when facing static coding tasks. Secondly, agentic tasks such as SWE-bench are inherently different in format and capabilities evaluated from non-agentic tasks such as BigCodeBench, with the former focusing on context and project understanding, while the latter stressing algorithm design and instruction following. Furthermore, we hypothesize that this effect might be largely influenced by the design of the agent scaffold used during training.
@@
Qwen2.5-Coder-7B - 01.81.8 $45.8$ 00.720.72 07.557.55 $1.82$ $17.58$
SWE-Gym-7B 491 $10.6$ $33.8$ 01.451.45 05.195.19 $2.51$ $12.16$
R2E-Gym-7B 3.3k $19.0$ 00.720.72 02.662.66 $2.54$ $16.90$
SWE-smith-7B 2.0k $15.2$ 00.000.00 $12.30$ $2.62$ 06.086.08
SWE-Play-mix-7B 704 $17.0$ 06.46.4 03.263.26 $30.69$ $2.95$ $14.19$
SWE-Play-general-7B 280 08.68.6 $22.6$ 02.542.54 07.557.55 $2.62$ $14.86$
SWE-Play-swe-7B 213 $11.0$ $10.6$ 00.720.72 04.804.80 $2.38$
@@
Qwen2.5-Coder-32B - 07.07.0 09.59.5 09.429.42 $24.81$ $2.31$
SWE-Gym-32B 491 $20.6$ $13.8$ 03.263.26 $10.34$ $2.51$
R2E-Gym-32B 3.3k $34.4$ 03.263.26 $15.04$ $3.30$
SWE-smith-32B 5.0k $40.2$ $13.77$ $42.15$ $2.62$
SWE-Play-mix-32B 704 $31.2$ 06.86.8 $18.12$ $44.24$ $3.64$